OpenStudy (anonymous):

I'm trying to find the limit when x->(-)infinity of (sqrt(x^6+3x^5+1)) / (7x^3+2x^5) what do i do with the sqrt in the numerator? is the answer infinity or sqrt(1)/7?

OpenStudy (anonymous):

\[\sqrt(x^6+3x^5+1)/(7x^3+2x^5)\]

OpenStudy (anonymous):

\[\lim_{x \rightarrow -\infty} \sqrt{x^6+3x^5+1}/(7x^3+2x^5)\]

OpenStudy (unklerhaukus):

\[\lim\limits_{x\rightarrow-∞} \frac{ \sqrt {{x^6+3x^5+1}}} {7x^3+2x^5 }\]

OpenStudy (anonymous):

does the \[\sqrt{x^6}\] turn into \[x^3\]?

OpenStudy (unklerhaukus):

in the limit that x is large and negative the important terms are the \[\sqrt {x^6}\text{, and the }2x^5\]

OpenStudy (anonymous):

but does the sqrt simplify to an x^3 or stay to the 6th power?

OpenStudy (anonymous):

x^3 / 2x^5 ->low degree/high degree -> lim=0

OpenStudy (anonymous):

or 6th power / 5th power -> high/low -> lim=infinity

OpenStudy (unklerhaukus):

yeah your right the \[√{x^6} \quad\text {would become}\quad x^3\] however x is negative so you cant do that

OpenStudy (anonymous):

hmmm, so does the answer not exist?

OpenStudy (unklerhaukus):

um i am still trying to figure it out

OpenStudy (anonymous):

okiedokie. take your time, this one has had me stumped for a long while...

OpenStudy (unklerhaukus):

i am just gonna use L'Hôpital's rule and see what happens

OpenStudy (zarkon):

factor out an x^6 on the top and x^5 on the bottom

OpenStudy (zarkon):

\[\frac{ \sqrt {{x^6+3x^5+1}}} {7x^3+2x^5 }\] \[=\frac{ \sqrt {{1+(3/x)+(1/x^6))}}} {x^5((7/x^2)+2)}\] \[=\frac{ |x^3|\sqrt {{x^6(1+(3/x)+(1/x^6))}}} {x^5((7/x^2)+2) }\] =...

OpenStudy (zarkon):

limit will be zero

OpenStudy (zarkon):

oops typo...

OpenStudy (zarkon):

\[=\frac{ |x^3|\sqrt {{1+(3/x)+(1/x^6)}}} {x^5((7/x^2)+2) }\]

OpenStudy (unklerhaukus):

do not attempt to use L'Hôpital's rule, you wil run out of paper Zarkon has correct answer

OpenStudy (anonymous):

oh, ok. So, when the x^6 gets pulled out it becomes an absolute value?

OpenStudy (zarkon):

\[\sqrt{x^6}=|x^3|\]

OpenStudy (anonymous):

thanks for going through the trouble @UnkleRhaukus!

OpenStudy (anonymous):

and you too Zarkon!

OpenStudy (unklerhaukus):

ah that is what to do if you have a negative index im glad learn something (thanks zarkon and thanks hicksonm